OpenStudy (pottersheep):

Calculus help please. Use implicit differentiation to show that any tangent line at a point P to a circle with center C is perpendicular to the radius CP. I'm not sure how to do this. Thanks!

OpenStudy (abb0t):

perpendicular: 

OpenStudy (pottersheep):

Yep

OpenStudy (abb0t):

Circle: \[x^2+y^2=1\]

OpenStudy (abb0t):

Use implicit differentiation on that.

OpenStudy (abb0t):

and solve for y'

OpenStudy (pottersheep):

2x + 2yy' = 0 yy' = -x y' = -x/y

OpenStudy (abb0t):

thus, the slope is x/y and you know that any two lines of a slope are perpendicular if say, fg = -1 Thus the tangent line and the radius are perpendicular.

OpenStudy (abb0t):

slope of CP = x/y

OpenStudy (pottersheep):

Sorry how do i know the slope of cp?

OpenStudy (abb0t):

your slope is ur tangent line.

OpenStudy (phi):

the question says, "with center C", so I would define C= (x0,y0) \[ (x-x_0)^2 + (y-y_0)^2 = r^2 \] implicitly differentiate: \[ 2 (x -x_0) dx + 2(y-y_0) dy = 0 \] \[ 2(y-y_0) dy = -2(x-x_0) dx \] \[ \frac{dy}{dx} = -\frac{(x-x_0)}{(y-y_0)} \] now find the slope of any radius at point (x,y): change in y over change in x : \[ \frac{y-y_0}{x-x_0} \] which is the negative reciprocal of the slope of the tangent

OpenStudy (pottersheep):

Thank you very much